# EDTformer: An Efficient Decoder Transformer for Visual Place Recognition

### v1

@@ Abstract
Visual place recognition (VPR) aims to determine the general geographical location of a query image by retrieving visually similar images from a large geo-tagged database. To obtain a global representation for each place image, most approaches typically focus on the aggregation of deep features extracted from a backbone through using current prominent architectures (e.g., CNNs, MLPs, pooling layer and transformer encoder), giving little attention to the transformer decoder. However, we argue that its strong capability in capturing contextual dependencies and generating accurate features holds considerable potential for the VPR task. To this end, we propose an Efficient Decoder Transformer (EDTformer) for feature aggregation, which consists of several stacked simplified decoder blocks followed by two linear layers to directly generate robust and discriminative global representations for VPR. Specifically, we do this by formulating deep features as the keys and values, as well as a set of independent learnable parameters as the queries. EDTformer can fully utilize the contextual information within deep features, then gradually decode and aggregate the effective features into the learnable queries to form the final global representations. Moreover, to provide powerful deep features for EDTformer and further facilitate the robustness, we use the foundation model DINOv2 as the backbone and propose a Low-Rank Parallel Adaptation (LoPA) method to enhance it, which can refine the intermediate features of the backbone progressively in a memory- and parameter-efficient way. As a result, our method not only outperforms single-stage VPR methods on multiple benchmark datasets, but also outperforms two-stage VPR methods which add a re-ranking with considerable cost. Code will be available at https://github.com/Tong-Jin01/EDTformer.
@@
Visual place recognition, feature aggregation, foundation models, parameter efficient transfer learning.
publicationid: pubid: $0$–$00$ 2021 IEEE
Figure 1: The performance comparison (Recall@1) on multiple benchmark datasets between our method and current state-of-the-art VPR methods, such as MixVPR [33], EngenPlaces [45], CricaVPR [24] and SelaVPR [25]. Our EDTformer consistently shows obvious advantages over other methods in diverse VPR scenarios, including viewpoint variations and condition changes (MSLS [39]), severe lighting changes (Tokyo24/7 [40]), various natural scenes (SPED [60]), place image modality changes (AmsterTime [42]) and varying weather conditions (SVOX [43]).
@@ I Introduction
VPR is generally addressed as a special image retrieval task [9]. Each place image is represented by a global feature, and then the nearest neighbor search is performed in the feature space to obtain the best-matched images of the query. The global features are typically derived from the aggregation of deep features, utilizing some techniques such as NetVLAD [12], GeM [14] pooling or their variants [16, 13, 44]. Unfortunately, such global features usually can not perform well in challenging VPR scenes. To further improve the robustness and retrieval accuracy, recent single-stage VPR research has focused on exploring novel methods to aggregate deep features. For instance, MixVPR [33] incorporates global relationships into each individual feature map through a stack of Feature-Mixer, which just consists of the multi-layer perceptrons (MLPs). CricaVPR [24] proposes a cross-image encoder to calculate the correlation between representations of multiple images in the same batch for robust global features. SALAD [46] redefines the soft assignment of local features in NetVLAD as an optimal transport problem, in which both CNNs and MLPs are fully utilized. BoQ [58] attempts to aggregate features by utilizing the stacked transformer encoders, attention mechanism and multiple sets of learnable queries. However, the transformer decoder structure has not been well explored in the VPR task, despite it being highly renowned in the fields of NLP [35] and semantic segmentation [56, 64, 65, 55] due to its strong capability in capturing contextual dependencies and generating accurate features. We argue that these characteristics of the decoder hold significant potential for addressing various challenges in VPR. To this end, we revisit the transformer decoder and propose an Efficient Decoder Transformer (EDTformer), which can fully utilize the contextual information within the deep features extracted from the backbone, then gradually decode and aggregate the crucial features to output robust and discriminative global representations for the place images.
@@
Moreover, as vision foundation models [20, 22, 21] have demonstrated a remarkable ability for feature extraction, recent VPR studies tend to use a pre-trained foundation model, e.g., DINOv2 [20], as the backbone to extract deep features from the input place images. AnyLoc [23] first applied DINOv2 for VPR without any fine-tuning, making it difficult to fully unleash the capability of DINOv2 for VPR. Then some research [58, 46] began to enhance the performance of DINOv2 in VPR through only fine-tuning the last few encoder blocks of DINOv2. Unfortunately, this process is accompanied by a large number of trainable parameters. Meanwhile, CricaVPR [24] and SelaVPR [25] attempted to apply parameter efficient transfer learning (PETL) methods [26, 49, 50] to adapt DINOv2 for VPR. Specifically, they froze DINOv2 and inserted some additional trainable adapters into each encoder block. However, this approach is efficient only in terms of parameters, not in training time and memory usage, as the gradient computation for the trainable parameters still requires backpropagation through the pre-trained backbone, as shown in Fig. 3 (b). This motivates us to develop a method that can fully strengthen the performance of vision foundation models in the VPR task in a memory- and parameter-efficient way.
@@
In this paper, we revisit the transformer decoder and propose a novel feature aggregation method EDTformer, which can finally generate robust and discriminative global representations for the place images. Without bells and whistles, to fully utilize the rich contextual information within the deep features, our EDTformer adopts a cascade of simplified decoder blocks, which only retain the attention (self-attention and cross-attention) layers and remove the feedforward network for improving the efficiency compared to the standard transformer decoder. By taking deep features as the keys and values, as well as utilizing a set of independent learnable parameters as queries, EDTformer can capture complex contextual relationships encoded in the deep features and progressively decode effective features into the learnable queries to achieve principal feature aggregation. Subsequently, the learnable queries which have contained crucial information are processed through two simple linear layers for dimensionality reduction and further adjustment, producing the final global representation. Meanwhile, we use the foundation model DINOv2 as the backbone in our architecture and develop a memory- and parameter-efficient Low-rank Parallel Adaptation (LoPA) method to adapt DINOv2 for the VPR task. Specifically, we freeze the entire backbone during training and design a tunable lightweight parallel network, which can progressively refine the intermediate features produced by each block of the backbone to output more powerful deep features, resulting in enhancing the performance of the whole model in VPR.
@@
(1) We revisit the transformer decoder and propose a novel feature aggregation method EDTformer, by leveraging several stacked simplified decoder blocks, linear projection and a set of learnable queries to fully decode the deep features and finally output a robust and discriminative global representation for global-retrieval-based VPR. This provides a new insight into how to apply the decoder structure for VPR.
@@ II-A Visual Place Recognition
In early VPR research, global features were developed by aggregating the hand-crafted local features (e.g., SIFT [28] and SURF [29, 30]), employing some classical aggregation algorithms, such as Bag of Words [31], Fisher Vector [63] and Vector of Locally Aggregated Descriptors (VLAD) [32]. With the great success of deep learning in compute vision tasks, current predominant VPR methods [12, 68, 48, 11, 4, 15, 16, 10, 47, 5, 13, 1, 44, 45, 23, 33, 58, 46, 24, 25] have preferred leveraging large amounts of deep features rather than hand-crafted local features to boost performance. Besides, the traditional aggregation algorithms are gradually replaced by trainable aggregation/pooling layers, such as NetVLAD [12] and GeM pooling [14]. NetVLAD utilizes a trainable generalized VLAD layer to aggregate deep features, typically tending to get a high-dimensional descriptor. In contrast, the Generalized Mean (GeM) pooling is a simple alternative that can produce compact global representations. However, such compact representations usually fall short of delivering satisfactory performance in challenging VPR scenarios. Hence, numerous works [15, 61, 44, 45, 33, 24, 46, 58] proposed various novel training strategies and aggregation algorithms to further improve the global representations for better retrieval accuracy. SFRS [15] proposed self-supervised image-to-region similarities to thoroughly exploit the potential of challenging positive images and their corresponding sub-regions for training a more robust VPR model. CosPlace [44] and EigenPlaces [45] framed the task of VPR training as a classification problem and utilized the San Francisco eXtra Large (SF-XL) dataset to train their VPR models effectively. MixVPR [33] proposed an all-MLP aggregation technique and trained the model utilizing the multi-similarity loss [37] with full supervision. CricaVPR [24] introduced a cross-image correlation-aware representation learning method to enhance the robustness of global features. SALAD [46] reinterpreted the soft assignment of local features in NetVLAD as an optimal transport problem, solving it by using the Sinkhorn algorithm [67]. BoQ [58] utilized multiple sets of queries to capture universal place-specific attributes by utilizing several stacked transformer encoders and attention mechanism. These studies only used global features and achieved a relatively good recognition performance.
@@
In addition, two-stage VPR methods [16, 10, 17, 62], which first search for the top-k candidate images in the database using global features and then re-ranks the candidates based on local features, are also an effective approach to further improve recognition performance. The re-ranking process usually either employs geometric consistency verification or leverages the learnable network to produce dense local features for similarity computation [17, 62, 25]. For instance, DHE-VPR [62] adopted a transformer-based deep homography estimation network to fit homography for fast and learnable geometric verification. SelaVPR [25] introduced a novel hybrid global-local adaptation method and directly used the dense local features in cross-matching for re-ranking. However, the re-ranking is generally time-consuming and demands a huge memory footprint as well as substantial storage space, particularly when dealing with large databases. These shortcomings restrict the applicability of two-stage methods in resource-constrained and large-scale VPR scenarios. Unlike two-stage methods improving performance at a substantial cost, our proposed method efficiently uses the decoder to aggregate deep features and directly produce highly robust and discriminative global representations for global-retrieval-based VPR.
@@ II-B Transformer Decoder Architecture
II-C Parameter Efficient Transfer Learning
The vision foundation models [20, 21, 22] trained on huge quantities of data, such as DINOv2 (trained on the large-scale curated LVD-142M dataset with the self-supervised strategy), possess the ability to extract powerful features from the input images and have achieved remarkable results on various downstream vision tasks. In order to reduce the number of training parameters while maintaining the strong capability of these foundation models, the PETL methods [26, 49, 27, 50] have been proposed and are widely applied in many areas, including the VPR field. For example, CricaVPR [24] enhanced the performance of DINOv2 in VPR by introducing multi-scale convolution adapters. SelaVPR [25] inserted the vanilla adapters into DINOv2 to achieve a hybrid global-local adaptation to produce both global and local features for the VPR task. By only tuning the built-in lightweight adapters without adjustment to the frozen pre-trained model, they are both efficient in terms of parameters. Nevertheless, the memory overhead during training is primarily dominated by the activations, not only parameters, which means that parameter efficiency is not equivalent to memory efficiency [53]. Unlike them, we design a Low-Rank Parallel Adaptation method to fully unleash the capability of the pre-trained vision foundation model without inserting any parameter into it, which is both parameter- and memory-efficient.
Figure 2: The pipeline to produce the robust and discriminative global representation for single-stage VPR. Firstly, the backbone with Low-rank Parallel Adaptation (i.e., DINOv2 with LoPA) is employed to extract powerful deep features of the input image. Next, the features undergo a linear transformation and are fed into each cross-attention layer as the keys and values. Additionally, we initial a set of learnable queries as the input queries for the first self-attention layer. After passing through $L$ decoder blocks, we can obtain the learned queries which have aggregated the crucial contextual features. Then we use two fully connected layers: one for dimensionality reduction and the other for further information aggregation by adjusting the number of queries. Finally, the output features are flattened and L2-normalized as the global representation of the place image.
@@ III Methodology
In this section, we first briefly introduce our backbone for feature extraction. Next, we propose the EDTformer to achieve feature aggregation efficiently. Then we introduce the LoPA to adapt the foundation model to provide more powerful features for EDTformer, improving the performance of the whole model in a memory- and parameter-efficient way. Finally, we describe the training strategy in our experiments.
@@ III-A Feature Extraction
In this work, we adopt the vision foundation DINOv2 as the backbone, which is based on Vision Transformer (ViT) [34]. To process the input image $I\in\mathbb{R}^{w\times h\times c}$, ViT first divides the image into $p\times p$ non-overlapping small patches, and then linearly projects them into $d$-dimensional tokens $x_{p}\in\mathbb{R}^{N\times d}$ ($N=p\times p$). Meanwhile, a learnable class token is prepended to $x_{p}$ to obtain $x_{0}=[x_{class},x_{p}]\in\mathbb{R}^{(N+1)\times d}$. Subsequently, position embeddings are added to $x_{p}$ to preserve positional information, resulting in $z_{0}$, which will be processed by a series of repeated transformer encoder blocks to extract features. A transformer encoder block mainly consists of the Multi-Head Attention (MHA) layer, the Multi-Layer Perceptron (MLP), and the LayerNormalization (LN), as shown in Fig. 3 (a). The token sequence $z_{l-1}$ goes through the transformer encoder block and produces $z_{l}$, which can be formulated as
@@ III-B EDTformer
It is worth noting that the latest work BoQ [58] and our EDTformer both utilize the learnable queries, but the architectures are completely different. BoQ uses multiple stacked encoders to deal with the deep features extracted from the backbone and applies separate attention mechanism, as well as multiple sets of independent queries to individually learn the output of each encoder. In contrast, we discard the encoder blocks and first attempt to directly utilize a purely decoder-based structure with only a set of queries to achieve the feature aggregation in VPR, which is simpler and more efficient.
Figure 3: Illustration of our Low-rank Parallel Adaptation method. (a) is a standard transformer encoder block in ViT. (b) is the popular PETL method based on adapters, which usually inserts the trainable adapters into the encoder blocks of the frozen backbone. (c) is the proposed LoPA method. We design it as a separate and lightweight network so that backpropagation does not pass through the frozen backbone to greatly reduce memory usage during training.

### v2

@@ Abstract
Visual place recognition (VPR) aims to determine the general geographical location of a query image by retrieving visually similar images from a large geo-tagged database. To obtain a global representation for each place image, most approaches typically focus on the aggregation of deep features extracted from a backbone through using current prominent architectures (e.g., CNNs, MLPs, pooling layer, and transformer encoder), giving little attention to the transformer decoder. However, we argue that its strong capability to capture contextual dependencies and generate accurate features holds considerable potential for the VPR task. To this end, we propose an Efficient Decoder Transformer (EDTformer) for feature aggregation, which consists of several stacked simplified decoder blocks followed by two linear layers to directly produce robust and discriminative global representations. Specifically, we do this by formulating deep features as the keys and values, as well as a set of learnable parameters as the queries. Our EDTformer can fully utilize the contextual information within deep features, then gradually decode and aggregate the effective features into the learnable queries to output the global representations. Moreover, to provide more powerful deep features for EDTformer and further facilitate the robustness, we use the foundation model DINOv2 as the backbone and propose a Low-rank Parallel Adaptation (LoPA) method to enhance its performance in VPR, which can refine the intermediate features of the backbone progressively in a memory- and parameter-efficient way. As a result, our method not only outperforms single-stage VPR methods on multiple benchmark datasets, but also outperforms two-stage VPR methods which add a re-ranking with considerable cost. Code will be available at https://github.com/Tong-Jin01/EDTformer.
@@
Visual place recognition, feature aggregation, foundation models, parameter-efficient transfer learning.
publicationid: pubid: Copyright ©2025 IEEE. Personal use of this material is permitted. However, permission to use this material for any other purposes must be obtained from the IEEE by sending an email to pubs-permissions@ieee.org.
[…]
@@ I Introduction
VPR is generally addressed as a special image retrieval task [18]. Each place image is represented by a global feature, and then the nearest neighbor search is performed in the feature space to obtain the best-matched images of the query. The global features are typically derived from the aggregation of deep features, utilizing some techniques such as NetVLAD [19], GeM [20] pooling or their variants [21, 22, 23]. Unfortunately, such global features usually can not perform well in challenging scenes. To get robust global features and thereby address the main challenges in VPR, recent single-stage VPR research has focused on exploring novel methods to achieve feature aggregation using some popular architectures. For instance, MixVPR [1] incorporates global relationships into each feature map through a stack of Feature-Mixer, which just consists of the multi-layer perceptrons (MLPs). CricaVPR [3] proposes a cross-image encoder to calculate the correlation between representations of multiple images in the same batch for robust global features. SALAD [24] redefines the soft assignment of local features in NetVLAD as an optimal transport problem, in which both CNNs and MLPs are fully utilized. BoQ [5] attempts to aggregate features by utilizing additional transformer encoders, attention mechanism, and multiple independent sets of learnable queries. However, the transformer decoder structure has not yet been well explored in the VPR task, despite it being highly renowned in the fields of NLP [25] and semantic segmentation [26, 27, 28, 29] due to its strong capability in capturing contextual dependencies and producing accurate features. We argue that these characteristics of the decoder hold significant potential for addressing various challenges in VPR. To this end, we revisit the transformer decoder and propose an Efficient Decoder Transformer (EDTformer), which can fully utilize the contextual information within deep features extracted from the backbone, then gradually decode and aggregate the crucial features to output robust and discriminative global representations for the place images.
@@
Moreover, as vision foundation models [30, 31, 32] have demonstrated a remarkable ability for feature extraction, recent VPR studies tend to use a pre-trained foundation model, e.g., DINOv2 [30], as the backbone to extract deep features from the input place images. AnyLoc [33] first applied DINOv2 for VPR without any fine-tuning, making it difficult to fully unleash the capability of DINOv2 for VPR. Then some research (e.g., BoQ [5] and SALAD [24]) began to enhance the performance of DINOv2 in VPR through only fine-tuning the last few encoder blocks of DINOv2. Unfortunately, this process is accompanied by a large number of trainable parameters. Concurrently, CricaVPR [3] and SelaVPR [4] attempted to apply parameter-efficient transfer learning (PETL) methods [34, 35, 36] to adapt DINOv2 to the VPR task. Specifically, they froze DINOv2 and inserted additional trainable adapters into each encoder block. However, this approach is efficient only in terms of parameters, not in training time and memory usage, as the gradient computation for the trainable parameters still requires backpropagation through (nearly the entire) pre-trained backbone [37, 38, 39], as shown in Fig. 3 (b). This motivates us to develop a method that can fully strengthen the performance of vision foundation models in the VPR task in a memory- and parameter-efficient way.
@@
In this paper, we revisit the transformer decoder and propose a novel feature aggregation method EDTformer, which can finally produce robust and discriminative global representations for the place images. Without bells and whistles, to fully leverage the rich contextual information within deep features, EDTformer employs a cascade of our simplified decoder blocks. These blocks only retain the attention (self-attention and cross-attention) layers and remove the feedforward network for higher efficiency compared to the standard transformer decoder. By taking deep features as the keys and values, as well as introducing a set of learnable parameters as queries, EDTformer can capture complex contextual relationships encoded in the deep features and progressively decode effective features into the learnable queries to achieve principal feature aggregation. Subsequently, the learnable queries that have contained crucial information are processed through two simple linear layers for dimensionality reduction and further adjustment, producing the global representations. Meanwhile, we use the foundation model DINOv2 as the backbone of our framework and develop a memory- and parameter-efficient Low-rank Parallel Adaptation (LoPA) method to adapt DINOv2 for the VPR task. Specifically, we freeze the whole backbone during training and design a tunable lightweight parallel network, which can progressively refine the intermediate features output by each encoder block of DINOv2 to provide more powerful deep features for EDTformer, thus enhancing the robustness of our entire model.
@@
(1) We revisit the transformer decoder and propose a novel feature aggregation method EDTformer, by leveraging several stacked simplified decoder blocks, linear projection, and a set of learnable queries to fully decode the deep features and finally output a robust and discriminative global representation for global-retrieval-based VPR. This provides a new insight into how to apply the decoder structure for VPR.
@@ II-A Visual Place Recognition
In early VPR research, global features were developed by aggregating the hand-crafted local features (e.g., SIFT [40] and SURF [41, 42]), employing some classical aggregation algorithms, such as Bag of Words [43], Fisher Vector [44] and Vector of Locally Aggregated Descriptors (VLAD) [45]. With the great success of deep learning in compute vision tasks, current predominant VPR methods [19, 46, 47, 48, 17, 49, 21, 50, 51, 52, 22, 11, 23, 2, 33, 1, 5, 24, 3, 4] have preferred leveraging large amounts of deep features rather than hand-crafted local features to boost performance. Besides, the traditional aggregation algorithms are gradually replaced by trainable aggregation/pooling layers, such as NetVLAD [19] and GeM pooling [20]. NetVLAD utilizes a trainable generalized VLAD layer to aggregate deep features, typically tending to get a high-dimensional descriptor. In contrast, the Generalized Mean (GeM) pooling is a simple alternative that can produce compact global representations. However, such compact representations usually fall short of delivering satisfactory performance in challenging VPR scenarios. Hence, numerous works [49, 53, 23, 2, 1, 3, 24, 5] proposed various novel training strategies and aggregation algorithms to further improve the global representations for better retrieval accuracy. SFRS [49] proposed self-supervised image-to-region similarities to thoroughly exploit the potential of challenging positive images and their corresponding sub-regions for training a more robust VPR model. CosPlace [23] and EigenPlaces [2] framed the task of VPR training as a classification problem and utilized the San Francisco eXtra Large (SF-XL) dataset to train their VPR models effectively. MixVPR [1] proposed an all-MLP aggregation technique and trained the model utilizing the multi-similarity loss [54] with full supervision. CricaVPR [3] introduced a cross-image correlation-aware representation learning method to enhance the robustness of global features. SALAD [24] reinterpreted the soft assignment of local features in NetVLAD as an optimal transport problem, solving it by using the Sinkhorn algorithm [55]. BoQ [5] utilized additional transformer encoders to further process deep features from the backbone, and then used multiple independent sets of queries to capture universal place-specific attributes from the output of each encoder through employing attention mechanism. These studies only used global features and achieved a relatively good performance.
@@
In addition, two-stage VPR methods [21, 50, 56, 57], which first search for the top-k candidate images in the database using global features and then re-ranks the candidates based on local features, are also an effective approach to further improve recognition performance. The re-ranking process usually either employs geometric consistency verification or leverages the learnable network to produce dense local features for similarity computation [56, 57, 4]. For instance, DHE-VPR [57] adopted a transformer-based deep homography estimation network to fit homography for fast and learnable geometric verification. SelaVPR [4] introduced a novel hybrid global-local adaptation method and directly used the dense local features in cross-matching for re-ranking. However, re-ranking is generally time-consuming and demands a huge memory footprint as well as substantial storage, particularly when dealing with large databases. These shortcomings restrict the applicability of two-stage methods in resource-constrained and large-scale VPR scenarios. Unlike two-stage methods improving performance at a substantial cost, our proposed method efficiently uses the decoder to aggregate deep features and directly produce highly robust and discriminative global representations, thus addressing the main challenges in VPR.
@@ II-B Transformer Decoder Architecture
II-C Parameter-Efficient Transfer Learning
The vision foundation models [30, 32, 31] trained on huge quantities of data, such as DINOv2 (trained on the large-scale curated LVD-142M dataset with the self-supervised strategy), possess the ability to extract powerful features from the input images and have achieved remarkable results on various downstream vision tasks. In order to reduce the number of training parameters while maintaining the strong capability of these foundation models, the PETL methods [34, 35, 63, 36] have been proposed and are widely applied in many areas, including the VPR field. For example, CricaVPR [3] enhanced the performance of DINOv2 in VPR by introducing multi-scale convolution adapters. SelaVPR [4] inserted the vanilla adapters into DINOv2 to achieve a hybrid global-local adaptation to produce both global and local features for the VPR task. By only tuning the built-in lightweight adapters without adjustment to the frozen pre-trained model, they are both efficient in terms of parameter. Nevertheless, the memory overhead during training is also dominated by the activations, not only parameters, which means that parameter efficiency is not equivalent to memory efficiency [37, 38, 39]. Unlike them, we design a Low-rank Parallel Adaptation method to fully unleash the capability of the pre-trained vision foundation model without inserting any parameter into it, which is both parameter- and memory-efficient.
Figure 2: Our pipeline to produce the robust and discriminative global representation for single-stage VPR. Firstly, the frozen backbone with Low-rank Parallel Adaptation (i.e., DINOv2 with LoPA) is employed to extract powerful deep features of the input image. Next, the features undergo a linear transformation and are fed into each cross-attention layer as the keys and values. Additionally, we initial a set of learnable queries as the input queries for the first self-attention layer. After passing through $L$ our simplified decoder blocks, we can obtain the learned queries which have aggregated the crucial contextual features for the VPR task. Then we use two fully connected layers: one for dimensionality reduction and the other for further information aggregation by adjusting the number of queries. Finally, the output features are flattened and L2-normalized as the global representation of the place image.
@@ III Methodology
In this section, we first briefly introduce our backbone for feature extraction. Next, we propose the EDTformer method and describe in detail how it achieves feature aggregation efficiently. Then we introduce our LoPA method to adapt the foundation model in a memory- and parameter-efficient way to provide more powerful features for EDTformer, thus enhancing the robustness of our entire model. Finally, we describe the training strategy used in our experiments.
@@ III-A Feature Extraction
In this work, we adopt the vision foundation DINOv2 as the backbone, which is based on Vision Transformer (ViT) [64]. To process the input image $I\in\mathbb{R}^{w\times h\times c}$, ViT first divides the image into $p\times p$ non-overlapping small patches, and then linearly projects them into $d$-dimensional tokens $x_{p}\in\mathbb{R}^{N\times d}$ ($N=p\times p$). Meanwhile, a learnable class token is prepended to $x_{p}$ to obtain $x_{0}=[x_{class},x_{p}]\in\mathbb{R}^{(N+1)\times d}$. Subsequently, position embeddings are added to $x_{0}$ to preserve positional information, resulting in $z_{0}$, which will be processed by a series of repeated transformer encoder blocks to extract features. A transformer encoder block mainly consists of the Multi-Head Attention (MHA) layer, the Multi-Layer Perceptron (MLP), and the LayerNormalization (LN), as shown in Fig. 3 (a). The token sequence $z_{l-1}$ goes through the transformer encoder block and produces $z_{l}$, which can be formulated as
@@ III-B EDTformer
It is worth noting that the latest work BoQ [5] and our EDTformer both utilize learnable queries during feature aggregation, but our architectures are different. BoQ uses several additional transformer encoders to further deal with the deep features extracted from the backbone and applies separate attention mechanism, as well as multiple independent sets of queries to individually learn the output of each encoder. In contrast, we first attempt to directly utilize a purely decoder-based structure with only a set of learnable queries to consistently decode and aggregate the effective information from deep features, thus obtaining a robust and discriminative global representation, which is simpler and more efficient.
Figure 3: Illustration of our Low-rank Parallel Adaptation method. (a) is a standard transformer encoder block in ViT. (b) is the popular PETL method based on adapters, which usually inserts the trainable adapters into the encoder blocks of the frozen backbone. (c) is our proposed LoPA method. The intermediate features from each encoder block of the frozen DINOv2 are sequentially fed to the corresponding adaptation function together with the output from the previous adaptation function.
During training, backpropagation does not pass through the frozen backbone, thus greatly reducing memory usage.